# Cooper quartets designing in multi-terminal superconducting devices

### v1

@@
Cooper quartets designing in multi-terminal superconducting devices
Luca Chirolli NEST, Istituto Nanoscienze - CNR, and Scuola Normale Superiore, I-56127 Pisa, Italy    Alessandro Braggio NEST, Istituto Nanoscienze - CNR, and Scuola Normale Superiore, I-56127 Pisa, Italy    Francesco Giazotto NEST, Istituto Nanoscienze - CNR, and Scuola Normale Superiore, I-56127 Pisa, Italy
@@
Quantum design of Cooper quartets in a double quantum dot system coupled to ordinary superconducting leads is presented as a novel platform for the study of an elusive many-body state of matter, that is at the basis of the phenomenon of charge-$4e$ superconductivity. A fundamentally novel, maximally correlated ground state, in the form of a superposition of vacuum $|0\rangle$ and four-electron state $|4e\rangle$, emerges as a narrow resonance and it is promoted by an attractive interdot interaction. A novel phenomenology in the dissipationless transport regime is elucidated, that yields typical flux quantization in units of $h/4e$ and manifests in non-local multi-terminal coherence and in two-Cooper pair transport properties mediated by the quartet ground state. The results open the way to the exploration of correlation effects and non-local coherence in hybrid superconducting devices, parity-protected quantum computing schemes and more generally, the work poses the basis for the design and simulation of novel correlated states of matter starting from ordinary ingredients available in a quantum solid state lab.
@@
The Cooper instability predicts a state of matter in which electrons pair up and form a condensate showing unique properties such as zero resistivity and perfect diamagnetism. Making a conceptual jump we can ask ourselves if more complicate electron aggregates such as Cooper quartets could form and condense in a novel state that gives rise to charge-$4e$ superconductivity. In spin 1/2 systems, Pauli’s exclusion principle forbids local multiparticle aggregates beyond pairs and a Bardeen-Cooper-Schrieffer superconductor constitutes a weakly correlated condensate of pairs in which quartet correlations, that are absent in the mean-field ground state, may emerge only as a residual of interaction among pairs. Purely four-fermion instabilities have been predicted in higher spin systems Wu (2005) or in systems showing an additional quantum number, such as $\alpha$-particle in nuclear physics Röpke et al. (1998). It has been suggested that a charge-$4e$ superconducting state may emerge in the fluctuating state, where Cooper pairs are loosely bound, and in systems displaying a two-component condensate Berg et al. (2009); Lee (2014); Fradkin et al. (2015), such as copper-oxides Li et al. (2007); Ding et al. (2008), kagome systems Lin and Nandkishore (2021); Chen et al. (2021); Zhou and Wang (2022); Li et al. (2023); Guguchia et al. (2023); Yu (2023), twisted bilayer graphene Jian et al. (2021); Fernandes and Fu (2021); Liu et al. (2023), iron-based superconductors Fernandes et al. (2014), or doped topological insulators Matano et al. (2016); Shen et al. (2017). Effective charge-$4e$ superconducting transport properties are obtained in networks of ordinary Josephson junctions, where transport of Cooper pairs is suppressed by destructive interference effects and only two-Cooper pair transport can survive, yielding a phenomenology similar to the one associated to the proper charge-$4e$ state Douçot and Vidal (2002); Douçot et al. (2003); Protopopov and Feigel’man (2004); Douçot et al. (2005); Rizzi et al. (2006).
@@
In this work, we change the perspective and resort to the art of quantum design to tailor the conditions for the onset of a Cooper quartet ground state. First of all, in order to fully study the emergence of many-body correlations we consider a minimal model of a double quantum dot, that allows us to accommodate a four-electron superposition state through an additional orbital binary quantum number beyond spin. In addition, we exploit Cooper pairing as already provided by nature through proximity coupling the double-dot system to a ordinary superconductor. Finally, in order to stabilize the quartet phase we suggest a mechanism based on attractive interdot interaction, that despite its unusual character has been experimentally realized through a second double quantum dot Hamo et al. (2016), or through a transmission line resonator Delbecq et al. (2013). The system manifests a quartet ground state
@@
in which the vacuum state $|0\rangle$ and the four-electron state $|4e\rangle$ appear in superposition of equal weights and that describes the formation of a coherent, albeit minimal, condensate, whose properties are dictated by the charge $4e$. We show that many-body interactions among the dots yield a finite purely quartet correlator,
@@
with $d_{i\sigma}$ fermionic dot annihilation operators, showing that a quartet scattering amplitude has a purely interacting origin. The correlator $Q$ resonantly peaks to its extreme values ±$\pm 1/2$ when the states $|\phi_{Q}^{\pm}\rangle$ are realized. A manifestation of the charge $4e$ nature of the ground state appears when ordinary superconducting leads are attached to the double-dot system and a charge-$4e$ Josephson effect appears. Upon closing the double-dot system on a superconducting loop threaded by a magnetic flux, the ground state current yields a flux quantization in units of $h/4e$. The quartet ground state delocalized in the two quantum dots shows a strong non-local coherent character and we identify a novel non-local phase response in a three-terminal setup that allows us to harvest the correlated nature of the ground state.
These findings open the way to interaction effects in superconductors, where non-local phase coherence can be establish in multi-terminal setups, and it can find application in parity-protected quantum computing schemes, such as those based on two-Cooper pair transport Blatter et al. (2001); Ioffe et al. (2002); Ioffe and Feigel’man (2002); Brooks et al. (2013); Douçot and Ioffe (2012). More generally, the results presented open the way to the design and simulation of novel phases of matter based on exotic electron complexes. Through manipulation of electron interactions in few body systems with tools and ingredients available in a solid-state quantum lab, it is possible to study the properties of otherwise inaccessible states of matter.
The model.—We consider the double quantum dot system, with the two dots labeled with $i=1,2$, tunnel coupled to a common superconducting lead $S$ at the left of the system, as schematized in Fig. 1a). We describe their general Hamiltonian $H_{0}$ through single-electron gate-tunable levels with energy $\epsilon_{i}$, an intradot generally repulsive Hubbard term with strength $U$, and an interdot density-density interaction with strength $W$, that we specifically allow to be negative, so to describe an attraction between the dots Hamo et al. (2016); Delbecq et al. (2013). The dot Hamiltonian then reads
@@
with $n_{i}=n_{i\uparrow}+n_{i\downarrow}$, $n_{i\sigma}=d^{\dagger}_{i\sigma}d_{i\sigma}$, and $d_{i\sigma}$ dot fermionic annihilation operators. We then describe tunnel coupling to the superconducting lead $S$ as a proximity effect, for which the quantum dots develop a finite pairing amplitude to form Cooper pairs, either locally on each dot or in a delocalized way, between the two dots, so that the pairing Hamiltonian reads
@@
This model has been thoroughly studied in the limit in which the scattering off the superconducting lead becomes elastic and integration away of the leads in the $\Delta\to\infty$ limit produces effective pairing amplitude, $\gamma_{S}=2\pi\nu_{F}t^{2}_{DS}$ and $\gamma^{\prime}_{S}=\frac{1}{2}\gamma_{S}e^{-\delta r/\xi}\sin(k_{F}\delta r)/(k_{F}\delta r)$, that coincides with the rates of local and crossed Andreev reflection, with $t_{DS}$ the dot-superconductor tunneling amplitude, assumed equal for the two dots, $\delta r$ the interdot distance, $k_{F}$ the Fermi momentum, and $\xi$ the coherence length of the superconductor. The full Hamiltonian is $H=H_{0}+H_{p}$ and we focus the analysis on the even parity sector, where the four electron operator $d_{1\downarrow}d_{1\uparrow}d_{2\downarrow}d_{2\uparrow}$ can have non-zero matrix element.
Quartet correlations emerge specifically when the resonance $04\epsilon+2U+4W=$ 0 occurs, for which the vacuum $|0\rangle$ and the four-fold occupied state $|4e\rangle$ become degenerate. To further select them as a host for a quartet ground state, we require all other states to have higher energy and, assuming a repulsive onsite interaction $U>0$, a relevant condition is met when the quantum dots experience an attractive density-density interaction that is stronger than the onsite repulsion, $-W>U>0$. At resonance, $\epsilon=-U/2-W>0$, and for $-2W>0$, $-U-W>0$, and $\gamma_{S},\gamma_{S}^{\prime}\ll|U+W|,U,|W|$, the vacuum and the four-electron states couple at first order with the higher energy states $|\phi_{2e}^{+}\rangle\equiv(|{\uparrow\downarrow}\rangle_{1}+|{\uparrow\downarrow}\rangle_{2})/\sqrt{2}$ and $|\phi_{2e}^{-}\rangle\equiv(|{\uparrow}\rangle_{1}|{\downarrow}\rangle_{2}-|{\downarrow}\rangle_{1}|{\uparrow}\rangle_{2})/\sqrt{2}$, as schematized in Fig. 1b) and detailed in  [^1] Supplemental Material containing details on the Hamiltonian and the correlator., and results in the low energy states
$|\bar{0}\rangle=|0\rangle+\frac{\sqrt{2}\gamma_{S}}{2W}|\phi^{+}_{2e}\rangle+\frac{\sqrt{2}\gamma^{\prime}_{S}}{U+W}|\phi^{-}_{2e}\rangle,$ (5)
$|\bar{4e}\rangle=|4e\rangle+\frac{\sqrt{2}\gamma_{S}}{2W}|\phi^{+}_{2e}\rangle-\frac{\sqrt{2}\gamma^{\prime}_{S}}{U+W}|\phi^{-}_{2e}\rangle.$ (6)
@@
with the coupling matrix element $\Gamma=-\frac{2(\gamma_{S}^{\prime})^{2}}{U+W}+\frac{\gamma_{S}^{2}}{W}$ and $\delta\epsilon$ a weak detuning from resonance, $\epsilon=-U/2-W+\delta\epsilon$.
Figure 2: a) Schematics of the three-terminal structure composed by two quantum dots, 1 and 2, coupled to a common superconducting lead $S_{0}$ at the left and to superconducting leads, $S_{1}$ and $S_{2}$, at the right. The latter are biased at phase $\varphi_{1}$ and $\varphi_{2}$ with respect to $S_{0}$, respectively. b) Contour plot of the ground state energy of the system in a) with respect to the phases $\varphi_{1}$ and $\varphi_{2}$, showing two minima at approximately $\varphi_{1}=\varphi_{2}\simeq\pm\pi/2$, for $\gamma_{S}^{\prime}=\gamma_{S,l}=\gamma_{S,r}=\gamma_{S}$, $\epsilon=-U/2-W$, $U/\gamma_{S}=10$ and $W/\gamma_{S}=-12$. c) Total current $I_{1}+I_{2}$ as a function of $\varphi_{1}=\varphi_{2}$ along the white line in b), shown for the case $\gamma_{S,l}=\gamma_{S,r}$ as in b) and for $\gamma_{S,l}=0$ for which the current is exactly $\pi$-periodic.
@@
The phase diagram of the system is shown in Fig. 1c), where we plot the ground state value of correlator $Q$ at resonance $\epsilon=-U/2-W$, for positive and negative values of $U$ and $W$. The correlator can take generic complex values such that $|Q|\leq 1/2$ and for real $\gamma_{S},\gamma_{S}^{\prime}$ it is real. In the region $U>W$ the ground state belongs to the subspace spanned by $\{|\phi^{-}_{Q}\rangle,|\phi^{-}_{2e}\rangle\}$ and the correlator evolves from $Q=-1/2$, when $U+W<0$ and the ground state has a strong $|\phi^{-}_{Q}\rangle$ quartet character, to $Q\sim(\gamma_{S}^{\prime})^{2}/(U+W)^{2}$ when $U+W>0$, where the ground state has mostly a two-electron state character $|\phi^{-}_{2e}\rangle$. In the region $U<W$ of the phase diagram Fig. 1c) the ground state belongs to the sector spanned by $\{|\phi^{+}_{Q}\rangle,|\phi^{+}_{2e}\rangle\}$, and for the fully attractive case $W,U<0$ a quartet ground state $|\phi^{+}_{Q}\rangle$ is achieved for which $Q=1/2$. Moving away from resonance, for $\epsilon=-U/2-W+\delta\epsilon$, the value of the correlator gives a Breit-Wigner resonance $Q\simeq-\Gamma/(2\sqrt{\Gamma^{2}+4\delta\epsilon^{2}})$ with linewidth $\Gamma$ for $U+W<0$, as shown in Fig. 1d), and away from resonance the quartet correlations rapidly decaying.

### v2

# Cooper quartets in interacting hybrid superconducting systems

@@
Abstract
@@
Cooper quartets represent exotic fermion aggregates describing correlated matter at the basis of charge-$4e$ superconductivity and offer a platform for studying four-body interactions, of interest for topologically protected quantum computing, nuclear matter simulations, and more general strongly correlated matter. Focusing on solid-state systems, we show how to quantum design Cooper quartets in a double-dot system coupled to ordinary superconducting leads through the introduction of an attractive interdot interaction. A fundamentally novel, maximally correlated double-dot ground state, in the form of a superposition of vacuum $|0\rangle$ and four-electron state $|4e\rangle$, emerges as a narrow resonance in a many-body quartet correlator that is accompanied by negligible pair correlations and features a rich phenomenology. The system represents an instance of correlated Andreev matter and the results open the way to the exploration of interaction effects in hybrid superconducting devices, and the study of novel correlated states of matter with ingredients available in a quantum solid-state lab.
I Introduction
@@
The Cooper instability predicts a state of matter in which electrons pair up and form a condensate showing unique properties such as zero resistivity and perfect diamagnetism. Making a conceptual jump we can ask ourselves if more complex electron aggregates such as Cooper quartets could form and condense, yielding charge-$4e$ superconductivity. Such a highly correlated fermion state belongs to a family of complex fermion states, that are typically the subject of intensive quantum simulations through quantum computing platforms Feynman (1982); Lloyd (1996); Abrams and Lloyd (1997); Aspuru-Guzik et al. (2005); Georgescu et al. (2014), and its isolation could prove useful for topological quantum computation Kitaev (2003), simulation of nuclear systems Röpke et al. (1998); Funaki et al. (2008); Schuck (2013) and quantum gravity Volovik (2024), and it could offer great insight in the study of four-body and many-body interactions Mizel and Lidar (2004); Peng et al. (2009); Dai et al. (2017); Kumar et al. (2020); Zhang et al. (2022). In spin 1/2 systems, Pauli’s exclusion principle forbids local multiparticle aggregates beyond pairs. Purely four-fermion instabilities have been predicted in higher spin systems Wu (2005), or in systems showing an additional quantum number, such as $\alpha$-particle in nuclear physics Röpke et al. (1998); Funaki et al. (2008); Schuck (2013). It has been suggested that a charge-$4e$ superconducting state may emerge in the fluctuating state of systems displaying a two-component condensate Babaev et al. (2004); Berg et al. (2009); Herland et al. (2010); Lee (2014); Fradkin et al. (2015), such as copper-oxides Li et al. (2007); Ding et al. (2008), kagome systems Lin and Nandkishore (2021); Chen et al. (2021); Zhou and Wang (2022); Li et al. (2023); Guguchia et al. (2023); Yu (2023), twisted bilayer graphene Jian et al. (2021); Fernandes and Fu (2021); Liu et al. (2023), iron-based superconductors Fernandes et al. (2014), or doped topological insulators Matano et al. (2016); Shen et al. (2017). In those systems, novel higher-order topological properties are emerging in fermion-quadrupling states and higher composites Grinenko et al. (2021); Babaev (2024); Maccari and Babaev (2022); Maccari et al. (2023). Effective charge-$4e$ superconducting transport properties are obtained in networks of ordinary Josephson junctions, where transport of individual Cooper pairs is suppressed by destructive interference effects and only two-Cooper pair transport can survive Douçot and Vidal (2002); Douçot et al. (2003); Protopopov and Feigel’man (2004); Douçot et al. (2005); Rizzi et al. (2006); Ciaccia et al. (2024), or in multiterminal Andreev bound state in the dissipative regime Cuevas and Pothier (2007); Freyn et al. (2011); Jonckheere et al. (2013); Feinberg et al. (2015); Cohen et al. (2018); Huang et al. (2022); Melo et al. (2022), yielding a phenomenology similar to the one associated to the proper charge-$4e$ state.
@@
In this work, we change the perspective and tailor the conditions for the onset of a Cooper quartet ground state. We consider a minimal model of a double quantum dot, that allows us to accommodate a four-electron superposition state through an additional orbital binary quantum number beyond spin. We exploit Cooper pairing through proximity coupling the double-dot system to a Bardeen-Cooper-Schrieffer (BCS) superconductor. We then add a crucial ingredient, that consists in an strong attractive interaction between the quantum dots and that allows us to stabilize the quartet superconducting phase. An attractive interdot interactions has be experimentally realized through a second double quantum dot Hamo et al. (2016), or through a transmission line resonator Delbecq et al. (2013), and it has been proposed to emerge by coupling to a flexural phonon in a suspended carbon nanotube Bhattacharya et al. (2021); Vigneau et al. (2022); Moser et al. (2013). The system manifests a quartet ground state
@@
in which the vacuum state $|0\rangle$ and the four-electron state $|4e\rangle$ appear in superposition of equal weights and that describes the formation of a coherent, albeit minimal, condensate, whose properties are dictated by the charge $4e$. We show that many-body interactions yield a finite quartet correlator,
@@
with $d_{i\sigma}$ fermionic dot annihilation operators, showing that pure quartet superconductivity violates the Wick theorem and has a genuine interacting origin. It is important to stress indeed that quartet correlations may be present in a BCS superconductor, but are subdominant with respect to pair correlations. The correlator $Q$ resonantly peaks to its extreme values ±$\pm 1/2$ when the states $|\phi_{Q}^{\pm}\rangle$ are realized, whereas the pair correlator is zero on these states. We study the Josephson currents by attaching extra BCS leads and single out the conditions for a two-Cooper pair current. More generally, the system realizes a correlated Andreev matter beyond the family of multi-terminal Josephson junctions Riwar et al. (2016); Teshler et al. (2023); Ohnmacht et al. (2024); Zalom et al. (2024), that is generated by interactions beyond the standard BCS mean field terms. The quartet ground state delocalized in the two quantum dots shows a strong non-local coherent character and we identify a novel non-local phase response in a three-terminal setup as a signature of the correlated nature of the ground state.
These findings open the way to the exploration of interaction-based superconductors and non-local phase coherence of multiterminal setups can be used to better investigate those systems Coraiola et al. (2023); Pankratova et al. (2020); Matsuo et al. (2022). Applications in parity-protected quantum computing schemes, such as those based on two-Cooper pair transport Blatter et al. (2001); Ioffe et al. (2002); Ioffe and Feigel’man (2002); Brooks et al. (2013); Douçot and Ioffe (2012); Brosco et al. (2024); Coppo et al. (2024), and more generally in the simulation of novel phases of matter constituted by exotic electron complexes can be envisioned using simple tools and ingredients available in a solid-state quantum lab.
II The model
We consider the double quantum dot system schematized in Fig. 1a), with the two dots labeled with $i=1,2$, tunnel coupled to a common superconducting lead $S$ at the left of the system. We describe their Hamiltonian $H_{0}$ through single-electron gate-tunable levels with energy $\epsilon_{i}$, an intradot (generally repulsive) Hubbard term with strength $U$, and an interdot density-density interaction with strength $W$. To theoretically investigate the rich phase diagram of the system we allow interactions to be attractive, and refer to possible underlaying mechanisms that have been theoretically and experimentally investigated Hamo et al. (2016); Delbecq et al. (2013); Bhattacharya et al. (2021); Vigneau et al. (2022); Moser et al. (2013). The double-dot Hamiltonian reads
@@ II The model
with $n_{i}=n_{i\uparrow}+n_{i\downarrow}$, $n_{i\sigma}=d^{\dagger}_{i\sigma}d_{i\sigma}$, and $d_{i\sigma}$ dot fermionic annihilation operators. We include tunnel coupling to the superconducting lead $S$ as a proximity effect, for which the quantum dots develop a finite pairing amplitude to form Cooper pairs, either locally on each dot or in a delocalized way, between the two dots, so that the pairing Hamiltonian reads
@@
This model has been thoroughly studied in the infinite superconducting gap limit ($\Delta\to\infty$), where scattering off the superconducting lead becomes elastic and integration away of the leads produces effective pairing amplitude, $\gamma_{S}=2\pi\nu_{F}t^{2}_{DS}$ and $\gamma^{\prime}_{S}=\frac{1}{2}\gamma_{S}e^{-\delta r/\xi}\sin(k_{F}\delta r)/(k_{F}\delta r)$. These coincide with the rates of local and crossed Andreev reflection, in which a Cooper pair can either tunnel to one of the two dots or split between the two dots, respectively, with $\nu_{F}$ the density of states at the Fermi energy, $t_{DS}$ the dot-superconductor tunneling amplitude, assumed equal for the two dots, $\delta r$ the interdot distance, $k_{F}$ the Fermi momentum, and $\xi$ the coherence length of the superconductor.
The full Hamiltonian is
$H=H_{0}+H_{p},$ (5)
and we focus the analysis on the even parity sector, where the four electron operator $d_{1\downarrow}d_{1\uparrow}d_{2\downarrow}d_{2\uparrow}$ can have non-zero matrix element. We further restrict ourselves to the subspace spanned by the following four states: the vacuum $|0\rangle$, the four electron state $|4e\rangle\equiv|{\uparrow\downarrow}\rangle_{1}|{\uparrow\downarrow}\rangle_{2}$, and the two doubly occupied states $|\phi_{2e}^{+}\rangle\equiv(|{\uparrow\downarrow}\rangle_{1}+|{\uparrow\downarrow}\rangle_{2})/\sqrt{2}$ and $|\phi_{2e}^{-}\rangle\equiv(|{\uparrow}\rangle_{1}|{\downarrow}\rangle_{2}-|{\downarrow}\rangle_{1}|{\uparrow}\rangle_{2})/\sqrt{2}$. In this subspace the Hamiltonian reads
$H_{e}=\left(\begin{array}[]{cccc}0&0&\sqrt{2}\gamma_{S}&\sqrt{2}\gamma^{\prime}_{S}\\ 0&4\epsilon+2U+4W&\sqrt{2}\gamma_{S}&-\sqrt{2}\gamma_{S}^{\prime}\\ \sqrt{2}\gamma_{S}&\sqrt{2}\gamma_{S}&2\epsilon+U&0\\ \sqrt{2}\gamma^{\prime}_{S}&-\sqrt{2}\gamma^{\prime}_{S}&0&2\epsilon+W\\ \end{array}\right).$ (6)
Quartet correlations emerge specifically around the resonance condition
$4\epsilon+2U+4W=0,$ (7)
for which the vacuum $|0\rangle$ and the four-fold occupied state $|4e\rangle$ become degenerate. In addition, the Hamiltonian decouples in two subspaces spanned by $\{|\phi^{+}_{Q}\rangle,|\phi^{+}_{2e}\rangle\}$ and $\{|\phi^{-}_{Q}\rangle,|\phi^{-}_{2e}\rangle\}$. To select the doublet $|0\rangle$ and $|4e\rangle$ as a host for a quartet ground state, we require all other states to have higher energy. By realistically assuming a repulsive onsite interaction $U>0$, a relevant condition is met when the quantum dots experience an attractive density-density interaction that is stronger than the onsite repulsion, $-W>U>0$. Close to resonance, $\epsilon=-U/2-W+\delta\epsilon$, with $\delta\epsilon$ a weak detuning from resonance, and for $-2W>0$, $-U-W>0$, and $\gamma_{S},\gamma_{S}^{\prime}\ll|U+W|,U,|W|$, the vacuum and the four-electron states couple at first order with the higher energy states $|\phi_{2e}^{+}\rangle$and $|\phi_{2e}^{-}\rangle$, as schematized in Fig. 1b), and result in the low energy states
$|\bar{0}\rangle=|0\rangle+\frac{\sqrt{2}\gamma_{S}}{2W-2\delta\epsilon}|\phi^{+}_{2e}\rangle+\frac{\sqrt{2}\gamma^{\prime}_{S}}{U+W-2\delta\epsilon}|\phi^{-}_{2e}\rangle,$ (8)
$|\bar{4e}\rangle=|4e\rangle+\frac{\sqrt{2}\gamma_{S}}{2W-2\delta\epsilon}|\phi^{+}_{2e}\rangle-\frac{\sqrt{2}\gamma^{\prime}_{S}}{U+W-2\delta\epsilon}|\phi^{-}_{2e}\rangle.$ (9)
@@
with the coupling matrix element
$\Gamma=-\frac{2(\gamma_{S}^{\prime})^{2}}{U+W}+\frac{\gamma_{S}^{2}}{W}.$ (11)
@@
The phase diagram of the system is shown in Fig. 1c), where we plot the ground state value of correlator $Q$ at resonance $\epsilon=-U/2-W$. For completeness, we present results for positive and negative values of $U$ and $W$. The correlator can take generic complex values satisfying $|Q|\leq 1/2$ and, for real $\gamma_{S},\gamma_{S}^{\prime}$, it is real (see Appendix A). In the region $U>W$ the ground state belongs to the subspace spanned by $\{|\phi^{-}_{Q}\rangle,|\phi^{-}_{2e}\rangle\}$ and the correlator evolves from $Q=-1/2$, when $U+W<0$ and the ground state has a strong $|\phi^{-}_{Q}\rangle$ quartet character, to $Q\sim(\gamma_{S}^{\prime})^{2}/(U+W)^{2}$ when $U+W>0$, where the ground state has mostly a two-electron state character $|\phi^{-}_{2e}\rangle$. In the region $U<W$ of the phase diagram Fig. 1c) the ground state belongs to the sector spanned by $\{|\phi^{+}_{Q}\rangle,|\phi^{+}_{2e}\rangle\}$, and for the fully attractive case $W,U<0$ a quartet ground state $|\phi^{+}_{Q}\rangle$ is achieved for which $Q=1/2$. Moving away from resonance, for $\epsilon=-U/2-W+\delta\epsilon$, the value of the correlator gives a Breit-Wigner resonance $Q\simeq-\Gamma/(2\sqrt{\Gamma^{2}+4\delta\epsilon^{2}})$ with linewidth $\Gamma$ for $U+W<0$, as shown in Fig. 1d), and away from resonance the quartet correlations rapidly decaying. Importantly, when the ground state belongs to the sector spanned by $\{|\bar{0}\rangle,|\bar{4e}\rangle\}$ the pair correlator is on order $\gamma_{S}/W$ or $\gamma_{S}^{\prime}/(U+W)$.
III Correlated Andreev matter
Figure 2: a) Schematics of the three-terminal structure composed by two quantum dots, 1 and 2, coupled to a common superconducting lead $S_{0}$ at the left and to superconducting leads, $S_{1}$ and $S_{2}$, at the right. The latter are biased at phases $\varphi_{1}$ and $\varphi_{2}$ with respect to $S_{0}$, respectively. b) Contour plot of the ground state energy of the system in a) with respect to the phases $\varphi_{1}$ and $\varphi_{2}$, showing two minima at approximately $\varphi_{1}=\varphi_{2}\simeq\pm\pi/2$, for $\gamma_{S}^{\prime}=\gamma_{S,l}=\gamma_{S,r}=\gamma_{S}$, $\epsilon=-U/2-W$, $U/\gamma_{S}=10$ and $W/\gamma_{S}=-12$. c) Plot of the ground state and the first excited states as a function of the phases $\varphi_{1}$ and $\varphi_{2}$, representing a correlated Andreev matter and realizing a quartet Andreev qubit. d) Total current $I_{1}+I_{2}$ as a function of $\varphi_{1}=\varphi_{2}=\varphi$ along the white line in panel c), shown for the case $\gamma_{S,l}=\gamma_{S,r}$, as in b) and c), and for $\gamma_{S,l}=0$ for which the current is exactly $\pi$-periodic.
It becomes natural at this point to search for a direct consequence of having a quartet ground state and a smoking gun observable that can witness its presence. The most intuitive property that a quartet ground state is expected to show is a quartet dissipationless current. However, since the leads are BCS superconductors, the fundamental carriers are Cooper-pairs and a Josephson current will in general flow through higher energy two-electron states, at second order in the rates $\gamma_{S},\gamma_{S}^{\prime}$, resulting in a standard $2\pi$-periodic current-phase relation, despite the fact that the double-dot system is in a quartet ground state. In particular, in a specular two-terminal configuration with an additional superconducting lead on the right side, coupled with same $\gamma_{S},\gamma_{S}^{\prime}$, the rates all acquire a factor $2\cos(\varphi/2)$ and the resulting current is $2\pi$ periodic, reflecting the fact that the system can load and discharge individual Cooper pairs via higher energy states through multiple processes involving local and non-local Andreev reflection.
Nevertheless, the system features a strongly correlated ground state and we expect that under proper conditions it can mediate a non-trivial two-Cooper pair current, that appears as a $\pi$-periodic Josephson current. For this to occur, a peculiar interplay between local and non-local Andreev processes needs to be arranged. More generally, the delocalized nature of the quartet ground state suggests that the system can manifest a correlated behavior in a multi-terminal configuration and a more general class of correlated multi-terminal Andreev matter can emerge.